# Building, Reusing, and Generalizing Abstract Representations from Concrete Sequences

### v1

@@ 5 Related Work
Previous modeling work on abstraction can be divided into two categories. The first category implements abstraction as searching for commonalities in explicitly symbolic systems. These are predominately cognitive models meant to explain human abstraction and transfer behavior, such as humans can quickly solve a new problem by finding its conceptual analogies with old problems. Lu et al. (2021) implemented abstraction and concept analogy as finding a common graph structure between a ground problem and an analogical problem in a semantic relational network that captures this property. Another example is the copycat project (Hofstadter & Mitchell, 1994) that models how people generalize rules shared amongst a couple of examples by searching for rules from a network of concepts. These works model abstraction on symbolic, explicit representations of knowledge graphs but do not address how the explicit knowledge graph itself arises from perceptual data. Our work proposes a mechanism for learning explicit abstraction from sequences of perceptual experience.
The second category implements abstraction in implicit connectionist systems. Works on meta-learning models and LLMs suggest their ability to generalize across contexts and solve problems in a way similar to humans (Wei et al., 2022; Binz & Schulz, 2023) in some tasks while failing short on other abstract reasoning tasks (Fleuret et al., 2011). Meanwhile, abstraction has been argued to be implicitly present in artificial neural networks (Yee, 2019; Kozma et al., 2018; Johnston & Fusi, 2023; Ito et al., 2022) and biological neural activities (Bernardi et al., 2020; Goudar et al., 2023). However, the abstractions neural networks learn are challenging to interpret (Fan et al., 2021). Relative to this approach, our work provides a reductionist cognitive model that explicitly specifies the minimal components needed for a model to learn interpretable abstract structure from sequences. Besides cognitive models, our work also differs from other approaches to sequence learning, such as probabilistic context-free grammar (PCFG) (Jelinek et al., 1992) or adaptor grammar (Johnson et al., 2006). Our generative mechanism offers a probabilistic, hierarchical sequence generation model relying on chunk-based recursive generation and inventory growth rather than formal grammar rules. HVM’s learning is also inventory-based instead of manipulating many substructures and rules.
@@ 6 Discussion
Our work has limitations. One is that variables are only proposed to be embedded between chunks and cannot come at the beginning or end of sequences, restricting the location of discoverable variables. Secondly, representation learned later in iteration depends on the earlier acquired representations. Furthermore, sometimes the variable structure can be arbitrarily nested depending on the learned representations up until that moment. Future work may look at optimizing the structure of the representation via refactorization features during learning. Finally, our work focused on the cognitive resemblance of HVM and has not explicitly optimized the model specifically for computational efficiency. Future work may look into optimizing run-time efficiencies of models that exploit hierarchical structures in data adapting to specific application proposes.
Our work opens up interesting future directions both in cognitive science and machine learning. Previously, grammar learning, chunk learning, and statistical/associative learning were studied in isolation as distinct aspects of sequence learning. We propose via HVM that the three can be treated as finding invariant recurring chunks from sequences, the means of it (statistical/associative learning), and the consequence of it (learning grammar-like structures). Our work suggests a normative origin of concrete and abstract chunk learning as a learning agent uncovering the underlying entities that constitute perceptual sequences. Future work can relate this model to the development of concept understanding during development. Meanwhile, the close tie between abstraction and generalization also urges future hypothesis-driven research regarding preconditions for learning abstract concepts to acquire more abstract representations. This can help to illuminate the emergence of abstract representation during excessive data exposure and training and its relation to the generalizability of machine learning models (Power et al., 2022; Miller et al., 2024).
7 Conclusion
We propose a hierarchical variable learning model (HVM) that builds up entities of recurring abstract and concrete patterns from perceptual sequences, utilizing chunking, finding commonality, and variable proposal chunking and abstraction to unearth independently recurring entities in sequences with nested hierarchical structures. We show the model’s resemblance to human sequence learning and transfer and highlight the relation between abstraction and generalization. Our work is relevant for cognitive science and the AI community in understanding memory in humans and machines.
@@ 8 Reproducibility Statement
Detailed information about the HVM algorithm, proof, generative model, test and experimental details and results can be found in the supplementary information section. The code used for the algorithm and experiments will be available as a comment to the reviewers and area chairs as a link to an anonymous repository as soon as the discussion forum for all submitted papers is open.
@@ A.3 Approximate Recognition Inverse
In practice, identifying the ground truth representation graph is much more complicated, as the representations learned by HVM can be mapped to a context-sensitive grammar A.6, learning the correct chunks that exactly match the generative model relates to the non identifiability problem that multiple grammars can generate the same set of observed data (Post, 1946; Greibach, 1968). The sequence length comparable to the ground truth is possible to obtain. However, the dataset creation method does not guarantee a single optimal way, this is usually referred to as the undecidability problem in grammar induction literature: inferring the exact grammar from positive examples is undecidable. Also is the case for deciding if two grammar classes are equivalent. Our formulation of the generative model can be shown to be equivalent to context-sensitive grammar A.6, which is more complex and undecidable than context-free grammar. No general algorithm can take any finite set of examples and produce a grammar that will generate all and only those examples. Because of undecidability, heuristic and approximate methods are often used in practice in the linguistics grammar induction literature and formal language theory. These methods aim to find solutions that are good enough rather than exact ones. As the optimality of the ground truth cannot be guaranteed due to the common source of this problem with grammar induction, our approach restricts and simplifies the structures of patterns that are feasible for the HVM. Since there is a theoretical limit to what can be achieved when inferring grammatical structures from data. Our approach to HVM uses approximate, heuristic, and probabilistic methods to tackle this problem by imposing parsimonious computational characteristics that have been pronouncedly observed in the cognitive science learning literature. Instead of comparing the model with the specific chunks of the ground truth, we use a set of evaluation measures to evaluate the model’s performance.
@@ A.7 Rate-Distortion Theory
Rate-distortion theory specifies that the best possible compression of a signal $X$ contains a lower bound on quality loss specified by the Rate-Distortion Function (R(D)): $R(D)=\inf_{Q}R_{Q}\text{s.t.}\mathbb{E}[d(X,\hat{X})]\leq D$ (Shannon, 1959; Cover & Thomas, 2012). According to the $RD$ theory, the minimum $R$ (the amount of compression) at which information can be transmitted over a communication channel for a given level of information loss D is specified by the Rate-Distortion Function $(R(D))$: $R(D)=\min_{p(\hat{x}|x):E[d(X,\hat{X})]\leq D}I(X;\hat{X})$. The mutual information quantifies how much information needs to be preserved during compression to achieve this fidelity, and the function minimizes this quantity while satisfying the distortion constraint. Written in another way,$R(D)=\inf_{Q}R_{Q}\text{ s.t.}E[d(X,\hat{X})]\leq D$. $Q$ represents the encoding function that maps from $x$ to $\hat{x}$. RD is agnostic to the choice of the distortion function, and the representation complexity, which measures the nestedness of the variables learned by HVM.

### v2

@@ 5 Related Work
Previous modeling work on abstraction can be divided into two categories. The first is cognitive models, which characterize abstraction as searching for commonalities in explicitly symbolic systems (Lu et al., 2021); these models apply to explain human problem solving, finding problem solutions based on its conceptual analogies with the previously-solved problems (Hofstadter & Mitchell, 1994). These models describe knowledge using symbolic, explicit graphs but do not address how the representation arises from perceptual data. Other models similar in taste include the recent wake-sleep Bayesian models such as DreamCoder Ellis et al. (2021). Our work proposes a mechanism for learning explicit abstraction from sequences. HVM starts learning with an empty dictionary and does not assume program primitives. Additionally, the inference/recognition in the HVM is sufficiently simple and can be simple without extensive machinery.
The second model category implements abstraction in implicit connectionist systems. Works on meta-learning models and LLMs suggest their ability to generalize across contexts and solve problems in a way similar to humans in some tasks (Wei et al., 2022; Binz & Schulz, 2023) while failing short on other abstract reasoning tasks (Fleuret et al., 2011). Meanwhile, abstraction has been argued to be implicitly present in artificial neural networks (Yee, 2019; Kozma et al., 2018; Johnston & Fusi, 2023; Ito et al., 2022) and biological neural activities (Bernardi et al., 2020; Goudar et al., 2023), albeit challenging to interpret (Fan et al., 2021). Relative to this approach, our work provides a reductionist cognitive model that explicitly specifies the minimal components needed to learn interpretable abstract structures.
Besides cognitive models, our work relates to and differentiates from other sequence learning approaches. Our generative mechanism relies on chunk-based recursive generation and inventory growth rather than formal grammar rules (Jelinek et al., 1992; Johnson et al., 2006), Zipfian distributions (Teh, 2006), abruptly changing statistics (Milan et al., ), or sequences from a stochastic chain Willems et al. (1995). We suggest that the perceptual sequence contains nested hierarchical structures, with chunks as a central generative force.
The hierarchical nature of HVM shares with models including topics models and Helmholtz Machine; however, it differs from topics models Blei et al. (2003); Teh et al. (2006) aiming at sequences rather than ‘bags of words’. The Helmholtz Machine also contains hierarchical generative and recognition components, but its focus is similarly not on sequences, and its sequential extension Hinton et al. (1995) did not work well partly because a lack of iterative processing that HVM employs. Finally, the goal of HVM as a computational cognitive model differs from neural models such as CSCG (George et al., 2021) which has an implementation level focus Marr (1982). However, we encourage future work to investigate the neural underpinnings of abstraction.
@@ 6 Discussion
Our work has limitations. One is that variables are only proposed to be embedded between chunks and cannot come at the beginning or end of sequences, restricting the location of discoverable variables. Secondly, representation learned later in iteration depends on those earlier acquired. However, this aspect aligns with the order and curriculum dependency of human learning. Children demand a simple-to-complex curriculum and utter more complex compounds later in their development (Friedmann et al., 2021; Snow, 1972; Fernald et al., 1989). There is a limit to HVM’s expressivity: it won’t be able to capture a pattern of two chunks being a variable number of chunks apart. Sometimes, the variable structure can be arbitrarily nested depending on the representations learned up until that moment. Future work may look at optimizing the structure during learning. Finally, our work focused on the cognitive resemblance of HVM and has yet to explicitly optimize the model for computational efficiency, leaving space for future work.
Our work opens up future directions both in cognitive science and machine learning. Previously, grammar learning (Chomsky, 2013), chunk learning (Gobet et al., 2001), and statistical/associative learning (Saffran et al., 1999) were studied as distinct characteristics of human sequence learning. We propose via HVM that the three components can be treated as finding invariant recurring chunks from sequences via statistical/associative learning, resulting in learning grammar-like structures. Our work suggests a normative origin of concrete and abstract chunk learning as a learning agent uncovers the underlying entities that constitute perceptual sequences. Future work can relate this model to developing concepts or compare the model with human behavior and LLMs on more nested datasets. Meanwhile, the close tie between abstraction and generalization also urges future hypothesis-driven research regarding preconditions for learning abstract concepts to acquire more complex representations. This can help to illuminate the emergence of abstract representation through excessive training and its relation to the generalizability of connectionist models (Power et al., 2022; Miller et al., 2024). Finally, our work can act in the tokenization role to fuse with transformers or SSM to extract interpretable representations from complex sequence data while enjoying the power of history-dependent sequence prediction.
@@ 7 Reproducibility Statement
Detailed information about the HVM algorithm, proof, generative model, test and experimental details and results can be found in the supplementary information section. The code used for the algorithm and experiments is available under this link.
@@ A.3 Approximate Recognition Inverse
In practice, identifying the ground truth representation graph is much more complicated, as the representations learned by HVM can be mapped to a context-sensitive grammar A.6, which is more complex and undecidable than context-free grammar. Learning the correct chunks that exactly match the generative model relates to the non-identifiability problem that multiple grammars can generate the same set of observed data (Post, 1946; Greibach, 1968) - inferring the exact grammar from positive examples is undecidable. The sequence length comparable to the ground truth is possible to obtain. Instead of comparing the model with the specific chunks of the ground truth, we use a set of evaluation measures to evaluate the model’s performance.
@@ A.7 Rate-Distortion Theory
Rate-distortion theory specifies that the best possible compression of a signal $X$ contains a lower bound on quality loss specified by the Rate-Distortion Function (R(D)): $R(D)=\inf_{Q}R_{Q}\text{s.t.}\mathbb{E}[d(X,\hat{X})]\leq D$ (Shannon, 1959; Cover & Thomas, 2012). According to the $RD$ theory, the minimum $R$ (the amount of compression) at which information can be transmitted over a communication channel for a given level of information loss D is specified by the Rate-Distortion Function $(R(D))$: $R(D)=\min_{p(\hat{x}|x):E[d(X,\hat{X})]\leq D}I(X;\hat{X})$. The mutual information quantifies how much information needs to be preserved during compression to achieve this fidelity, and the function minimizes this quantity while satisfying the distortion constraint. Written in another way,$R(D)=\inf_{Q}R_{Q}\text{ s.t.}E[d(X,\hat{X})]\leq D$. $Q$ represents the encoding function that maps from $x$ to $\hat{x}$. RD is agnostic to the choice of the distortion function, and in this case we use the representation complexity, which measures the nestedness of the variables learned by HVM.